how can i solve for "a"? (5a-3(4-a))/2=5
Answer by rfer(16322)

(5a-12+3a)/2=5
(8a-12)/2=5
8a-12=5*2
8a-12=10
8a=10+12
8a=24
a=24/8
a=3